OpenStudy (anonymous):

how to put x + 4y = -8 into slope intercept form HELP PLEASE!!!!

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Do you know what a slope intercept form is?

OpenStudy (anonymous):

y=mx + b

OpenStudy (anonymous):

do you know what form x + 4y = -8 is in?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

that form is called the standard form.

OpenStudy (anonymous):

you told me slope intercept form was y= mx + b right?

OpenStudy (anonymous):

I will show you step by step

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

as you can see, x and y are both on left side. However, slope intercept form says y is the only one of the left. right?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

so then you want to separate y by moving x to the right side.

OpenStudy (anonymous):